Let g(x) =3x + 5 and f(x) = x2 + 2x – 7 . Find f(g(x)).
Answer #1

g (u) = 34+5 #C)= +24-7 f69(a)) = f(3x+5) = (3x+5) 2 (34+5)-7 să+ 304+25 +60+10-7 2/ 93 +36x+28 (Answer).